Estimate the approximate valence electron of boron and oxygen?

The shielding of electrons gives rise to an effective nuclear charge, Zeff which explains why boron is larger than oxygen. Estimate the approximate Zeff felt by valence electron of boron and oxygen, respectively?

A. +5 and +8
B. +3 and +6
C. +5 and +6
D. +3 and +8
E. +1 and +4
Answer #1
Concepts and reason

Effective Nuclear Charge

The effective nuclear charge is defined as the net attractive interaction between the positive charge of nuclear protons and the negative charge of valence electrons. It is always less than the actual nuclear charge of the atoms.

Shielding Effect

The shielding effect is the diminishing of attractive interaction between the nuclear protons and the valence electrons by inner electrons.

Fundamentals

Mathematical equation for the effective nuclear charge:

Zer=Z-S

Where,

The atomic number of the atom is Z.

The nonvalenced electrons or core electrons is S.

Boron
(Valence shell)
> n = 2
(core or Inner shell)
n = 1
ze-
- Nucleus

The atomic symbol for boron atom is.

The atomic number of boron is 5. (Z=5)

The ground state electronic configuration of boron atom is 1s? 2s²2p
corevalence

The nonvalence electrons or core electrons of boron are 2. (S=2)

The valence electrons of boron atom are 3.

Zer = Z-S
Substitute 5 for Z and 2 for S.
Zer = 5-2
Zer = 3
Zor = +3

oxygen
(Valence shell)
> n = 2
(core or Inner shell)
n
=
1
– Nucleus

The atomic symbol for oxygen atom is.

The atomic number of oxygen is 8. (Z=8)

The ground state electronic configuration of oxygen atom is 182 2s²2p
corevalence

The nonvalence electrons or core electrons of oxygen are 2. (S=2)

The valence electrons of oxygen atom are 6.

Zer = Z-S
Substitute 8 for Z and 2 for S.
Zer = 8-2
Zer = 6
Zoo = +6

Ans:

The approximate felt by a valence electron of boron and oxygen atoms are +3 and +6, respectively.